find the volume of the region between the cylinder z=y^2 and the xy plane that is bounded by the plane x=1 x=0 y=-1 y=1
1
Expert's answer
2021-07-28T17:52:35-0400

Region projection onto the Oxy plane:



Then

"V = \int {\int\limits_D {z(x,y)dxdy} = \int\limits_0^1 {dx} \int\limits_{ - 1}^1 {{y^2}} } dy = \frac{1}{3}\int\limits_0^1 {\left. {{y^3}} \right|_{ - 1}^1dx} = \frac{1}{3}\int\limits_0^1 {(1 + 1)dx = } \frac{2}{3}\left. x \right|_0^1 = \frac{2}{3}"

Answer: "V=\frac{2}{3}"